What is the angular velocity of a wheel that has a period of rotation of 2.5 seconds?
1
Expert's answer
2019-04-17T11:10:21-0400

The velocity of a wheel is


"\omega=\frac{2\pi}{T}"

"\omega=\frac{2\pi}{2.5}=2.51 \frac{rad}{s}"